OpenStudy (anonymous):

Solve. Show all steps used to solve. 2(n - 3) + 5 = 3(n - 1) Solve. Show all steps used to solve. 2(n - 3) + 5 = 3(n - 1) @Mathematics

OpenStudy (aroub):

first distribute.. 2n-6+5=3n-3 second add like terms.. 2n-1=3n-3 third put the variables alone... -3n+2n=-3+1 fourth add like terms again... -n=-2 finally just remove the negative... n=2 :D